OpenStudy (anonymous):

Find an equation for the nth term of the sequence. -4, -16, -64, -256, ...

OpenStudy (solomonzelman):

The common ratio is \(\large\color{blue}{a_n \div a_{n-1} }\)

OpenStudy (solomonzelman):

\(\large\color{blue}{a_n=a_{1} \times r^{n-1}}\) is the form

OpenStudy (campbell_st):

well each term is being multiplied by 4 so the sequence is geometric... and a term in a geometric series is \[a_{n} = a \times r^{n - 1}\] you know a = -4 and n = 4 so just substitute them

OpenStudy (solomonzelman):

No, n is not 4, he is asking for any nth term in general, as I see.

OpenStudy (anonymous):

an=-4*4^n-1

OpenStudy (solomonzelman):

4^(n-1) is not the same as 4^n-1

OpenStudy (anonymous):

oh, ok thanks

OpenStudy (solomonzelman):

yeah, just marking that, np

OpenStudy (anonymous):

making a new question, can you guys hop over?